OpenStudy (jensine):

how to solve the equation x^3-x^2-30x=0

OpenStudy (agreene):

x^3-x^2-30x=0 divide both sides by x, which will yield: x=0 or x^2-x-30=0 solve the x^2-x-30=0 for the other two solutions to the trinomial (use quadratic/completing the square).